# FactoFormer: Factorized Hyperspectral Transformers with Self-Supervised Pre-Training

### 11 Dec 2023 v2

@@ Abstract
Hyperspectral images (HSIs) contain rich spectral and spatial information. Motivated by the success of transformers in the field of natural language processing and computer vision where they have shown the ability to learn long range dependencies within input data, recent research has focused on using transformers for HSIs. However, current state-of-the-art hyperspectral transformers only tokenize the input HSI sample along the spectral dimension, resulting in the under-utilization of spatial information. Moreover, transformers are known to be data-hungry and their performance relies heavily on large-scale pre-training, which is challenging due to limited annotated hyperspectral data. Therefore, the full potential of HSI transformers has not been fully realized. To overcome these limitations, we propose a novel factorized spectral-spatial transformer that incorporates factorized self-supervised pre-training procedures, leading to significant improvements in performance. The factorization of the inputs allows the spectral and spatial transformers to better capture the interactions within the hyperspectral data cubes. Inspired by masked image modeling pre-training, we also devise efficient masking strategies for pre-training each of the spectral and spatial transformers. We conduct experiments on six publicly available datasets for HSI classification task and demonstrate that our model achieves state-of-the-art performance in all the datasets. The code for our model will be made available at https://github.com/csiro-robotics/FactoFormer.
@@
Hyperspectral image, transformer network, factorized, self-supervised pre-training.
@@ I Introduction
However, previous methods  [13, 14] mainly applied pure Vision Transformer (ViT) architecture for spectral encoding, using a simple yet effective spectral group tokenization strategy, but they have not fully exploited the spatial and spectral interactions present in hyperspectral image cubes. Furthermore, the limitation of annotated data is a significant challenge when training ViTs, as they rely heavily on large-scale training data. However, most of the available hyperspectral datasets contain limited labeled proportions, such as the Indian Pines, University of Pavia, and Houston 2013 datasets, which possess only 48.75%, 17.43%, and 2.26% of annotated data, respectively. Recently, Masked Image Modeling (MIM) [16, 17] has been introduced as a simple yet powerful self-supervised pre-training framework for Vision Transformers (ViT) without the use of labels. However, the MIM pre-training strategies often have been limited to a single modality (RGB).
@@
The proposed factorized architecture not only increases the efficiency and scalability of modeling spectral and spatial interactions but also enables factorized self-supervised pre-training. To mitigate the lack of large labeled data, we design novel masking strategies to effectively pre-train spectral and spatial transformers individually, leveraging a vast amount of unlabeled hyperspectral data. A random proportion of the input tokens are masked, and only the visible tokens are passed through spectral and spatial transformers, respectively. The latent space representations from each transformer, combined with the masked tokens, are then used to reconstruct the original information for the masked regions.
@@
  • We propose novel self-supervised pre-training strategies for the factorized spatial and spectral transformers, which allows us to better utilize the large proportion of unlabeled data in the public datasets.
@@ II-C Transformers for Hyperspectral Images
Transformer-based networks have revolutionized the entire natural language processing domain with a significant performance boost. Inspired by this, vision transformer (ViT) [27] was introduced later, where images are treated as a sequence of patches. This model achieved state-of-the-art performance for image classification with large-scale pre-training. Transformers have proven to be the ideal architecture for capturing long range dependencies through its multi-head self attention mechanism which processes input as a sequence of tokens. Considering the large number of spectral bands in HSI, which has high correlation and granular information, extracting these global dependencies are crucial for learning better features. Therefore, utilizing transformers and adapting it for learning spatial-spectral features in HSI would lead to better performance. However, there are only few works exploring the transformer architectures for HSI and its full potential in the domain is yet to be explored. Sun et al. [15] proposed a hybrid network that utilizes both CNNs and transformers. Initially, 3D and 2D CNNs are used to extract features, and then the extracted features are tokenized by Gaussian weighted tokenizer and sent to the transformer encoder. In this method, the main feature extraction is based on CNNs, where a single-layer transformer is utilized at the last layer to learn relationships between high-level semantic features. Mei et al. [28] proposed a group-aware hierarchical transformer, which has three stages where each stage is composed of grouped pixel embedding and transformer layers. Here, the proposed grouped pixel embedding comprises convolutions to learn local features, while the transformer is utilized to learn global dependencies. These hybrid approaches perform the initial feature extraction using convolutions where the transformers are used in order to find the global dependency of extracted local features. They did not explore how transformers can be directly used with HSI to learn spectral-spatial feature interactions without convolutions. SpectralFormer [13] utilized the ViT architecture and proposed a pure transformer-based network for HSI with an additional skip connection mechanism to fuse information between transformer layers. Here, the input is patched along the spectra by grouping neighboring bands. Even though the model performed well, the input samples are patched only along the spectral dimension, where the spatial information in HSIs is not fully exploited. Also, the model is trained in a fully supervised manner using limited training samples, which degrades the performance of transformers due to the lack of annotated data. Recently, MAEST [14] utilizes a similar transformer architecture as SpectralFormer[13] and added a self-supervised pre-training framework for only spectral dimension, which is briefly discussed in the next section. Peng et al. [29] proposed a spatial-spectral transformer architecture with cross-attention for HSI. Here, the input is tokenized in spatial and spectral dimensions and fused with a spatial-spectral cross-attention layer in the transformer block. In this method, principal component analysis is used initially to reduce the dimensionality; hence the model is not exposed to full spectral ranges which may cause the loss of fine-grained information. Also, the approach is fully supervised and trained with limited labeled data and the additional convolution layers used prior to tokenization adds more complexity to the model.
@@ II-D Self-Supervised Learning
In this section we have explored and analyzed related works that closely align with our proposed method. We propose FactoFormer: a factorized transformer with self-supervised pre-training to learn rich spectral-spatial feature representations from hyperspectral data. The spectral-spatial factorization is proposed in [25] as well, but the approach is fully supervised. Also, the input is pre-processed with PCA in order to reduce the spectral dimension, potentially leading to a loss of fine-grained information. The input is further processed with additional convolutions prior to sending it to the transformer. In contrast, FactoFormer adopts a different approach. We tokenize the raw input without any preprocessing and directly feed it to the transformers. This approach enables the model to learn more fine-grained spatial-spectral correlations present in the hyperspectral data, making it well-suited for the task at hand.In terms of self-supervised pre-training both MAEST [14] and proposed FactoFormer utilized the unlabelled data to pre-train the network. In MAEST [14], the input is patched only along the spectra by grouping neighbouring bands, whereas in FactoFormer we propose a factorized pre-training approach which enables learning rich spectral and spatial features separately during pre-training and jointly fine-tuning to learn the higher-order spectral-spatial features. Also compared to MAEST [14] our proposed method is computationally more efficient due to the simple linear layer decoder used to reconstruct only masked tokens whereas in MAEST [14] a heavy transformer-based decoder is used to reconstruct the full input.
@@ IV Methodology
In this section, we first give an overview of our proposed method FactoFormer. Next, we discuss the proposed novel factorized transformer architecture. Finally, we explain the self-supervised pre-training strategy used in our method. The overall architecture of the model is shown in Figure 1.
@@ IV-A FactoFormer
Our proposed FactoFormer learns spectral and spatial representations from hyperspectral data using factorized transformers. We process the entire input cube using both spatial and spectral transformers in order to capture the intrinsic spatial-spectral correlations. But the way we tokenize and compute self-attention to learn salient spatial and spectral information discriminates against one another. Further, with the fusion mechanism employed at the end, we capture more fine-grained higher-order spatial-spectral correlations. The factorized architecture enables factorized pre-training, where both transformers can be treated separately and pre-trained using specific self-supervised pre-text tasks.
@@
Figure 2: An overview of the proposed pre-training network for spectral transformer in FactoFormer.
IV-B Self-Supervised Pre-training in FactoFormer
Inspired by Masked Image Modeling (MIM), we design a new pre-training strategy for both spectral and spatial transformers. We utilized the large proportion of unlabeled data from the same dataset for self-supervised pre-training. As these unlabeled samples belong to the same image, environmental conditions, instrumental noise, and other affecting factors remain constant during pre-training. They are helpful for learning better feature representations. Both transformers are pre-trained separately with spatially consistent and spectrally consistent masking strategies to enhance their underlying feature representations.
Pre-training of Spectral Transformer: We apply patching to the input sample in the spectral dimension, followed by linear embedding and positional encoding as described in Eq. 6. Subsequently, we randomly mask tokens from the input sample, passing only the visible tokens through the spectral transformer. The latent space representations from the spectral transformer, combined with the masked tokens, are then used to reconstruct the original information for the masked regions. For this purpose, we utilize a simple linear layer as the decoder, in contrast to the heavy transformer-based decoder used in MAEST [14]. Finally, we use Mean Squared Error (MSE) as the loss function. As shown in Figure 2, masks are spatially consistent. Each token corresponds to the complete spatial information of a particular band in the input sample, allowing the model to learn better representations along the spectral dimension from spectrally patched tokens.
Pre-training of Spatial Transformer: We apply patching to the input in the spatial dimension and pass them through the spatial transformer along with positional embeddings as outlined in Eq. 8. A random proportion of the input tokens are masked, and only the visible tokens are passed through the spatial transformer. Similar to the spectral transformer, we utilize a simple linear layer as the decoder to reconstruct the masked regions and use MSE as the loss function. Notably, the spectrally consistent masks enable the transformer to learn meaningful representations from spatially patched tokens, as depicted in Figure 3.
Figure 3: An overview of the proposed pre-training network for spatial transformer in FactoFormer.
End-to-End Fine-tuning: We use the pre-trained weights of the spectral and spatial transformers to initialize FactoFormer. Then we fine-tune the network end-to-end in order to perform classification. Our proposed factorization technique enables FactoFormer to exploit the pre-trained weights and learn rich spectral and spatial feature representations jointly without imposing significant complexity to the network.
@@ V-A Datasets
We use six popular hyperspectral datasets to evaluate the performance of the proposed model and to compare it with other state-of-the-art methods. These datasets include Indian Pines [34], University of Pavia [35], and Houston 2013 [36], as well as the Wuhan UAV-borne hyperspectral image (WHU-Hi) dataset [37], which consists of three individual datasets: WHU-Hi-LongKou, WHU-Hi-HanChuan, and WHU-Hi-HongHu. The details on number of samples used for pre-training, fine-tuning and testing are shown in Table I.
@@
TABLE I: Details of Data used for Pre-training, Fine-tuning and Testing
@@ V-B Implementation Details
Network Architectures: We use a transformer architecture with five layers and four attention heads for both spectral and spatial transformers.We set the embedding size for spectral and spatial transformers as 32 and 64, respectively. Table II outlines the details of the network architectures. The same architectures were preserved during pre-training as well.
@@
Pre-training: We employ the Adam optimizer with a linear learning rate scheduler and train for 200 epochs. The batch size is set to 32 with an initial learning rate of $5e^{-4}$ and decayed by a factor of 0.9 for every 20 epochs. The pre-training setting is detailed in Table IV.
Fine-tuning: For fine-tuning, the pre-trained weights are used to initialize both spectral and spatial transformers and fine-tuned end-to-end. A learning rate search is performed, and the best for each dataset is reported. We use the Adam optimizer with a linear learning rate scheduler and set the batch size to 32. For both the Indian Pines and University of Pavia datasets, the fine-tuning model converges after 80 epochs, while for Houston and WHU-Hi datasets, it converges in just 40 epochs. The Finetuning setup is detailed in Table IV.
TABLE III: Pre-training Setting
@@ V-C State-of-the-art Methods
Transformer Based Networks: For transformer-based networks, we used the ViT architecture [27], SpectralFormer [13], and MAEST [14]. The ViT uses five layers, four attention heads, an embedding size of 64, and an MLP with a hidden dimension of 8. SpectralFormer[13] uses the same settings as the ViT with cross-layer adaptive fusion (CAF). CAF employs a skip connection mechanism to integrate the features of skipping encoder blocks. MAEST[14] is a self-supervised approach that utilizes Masked Autoencoders (MAE) to train an encoder-decoder network that reconstructs masked patches. Then a classification encoder with similar network architecture as SpectralFormer[13] is initialized with the pre-trained encoder weights and fine-tuned to perform classification.
@@ VI-A Comparison with the State-of-the-Art
When comparing FactoFormer with MAEST[14], which is the current state-of-the-art method using a transformer with a self-supervised pre-training approach, our method outperforms it with 7.15%, 4.13%, and 0.58% higher OA on Indian Pines, University of Pavia and Houston 2013 datasets respectively. It shows that our novel factorized approach with self-supervised pre-training improves performance significantly. Also, interestingly our models converge with less than 100 epochs during fine-tuning on all three datasets, whereas, in MAEST[14], the models are fine-tuned for more than 200 epochs. It again highlights that the factorized pre-training in FactoFormer learns rich feature representations, which enables fast convergence during fine-tuning. Also, FactoFormer is computationally more efficient in terms of the number of model parameters and run-time, which is briefly discussed in Section VI-C under computational efficiency.
Compared to SpectralFormer [13], which is a fully supervised transformer-based approach, FactoFormer outperforms it by 9.54% on Indian Pines, 4.12% on University of Pavia and 1.12% on Houston 2013 datasets highlighting the importance of factorization and pre-training. ViT[27] shows the least performance of all the transformer methods, as it is applied directly as in images for spectral tokens in HSI without adapting it to suit the structure of hyperspectral data. In general, SpectralFormer[13], MAEST[14], and FactoFormer perform better than all other classic backbone networks and conventional classifiers in all three datasets, demonstrating the effectiveness of transformer-based approaches for HSI. 2-D CNN[9] performs better compared to 1-D CNN[8], RNN[41] and miniGCN[42] on all three datasets. When comparing 2-D CNN [9], FactoFormer achieves 15.41%, 9.14%, and 5.41% higher accuracy on Indian Pines, University of Pavia and Houston 2013 datasets, respectively. This significant performance gap exhibits the effectiveness of FactoFormer over CNN-based approaches.

### 04 Jan 2024 v4

# FactoFormer: Factorized Hyperspectral Transformers with Self-Supervised Pretraining

@@ Abstract
Hyperspectral images (HSIs) contain rich spectral and spatial information. Motivated by the success of transformers in the field of natural language processing and computer vision where they have shown the ability to learn long range dependencies within input data, recent research has focused on using transformers for HSIs. However, current state-of-the-art hyperspectral transformers only tokenize the input HSI sample along the spectral dimension, resulting in the under-utilization of spatial information. Moreover, transformers are known to be data-hungry and their performance relies heavily on large-scale pretraining, which is challenging due to limited annotated hyperspectral data. Therefore, the full potential of HSI transformers has not been fully realized. To overcome these limitations, we propose a novel factorized spectral-spatial transformer that incorporates factorized self-supervised pretraining procedures, leading to significant improvements in performance. The factorization of the inputs allows the spectral and spatial transformers to better capture the interactions within the hyperspectral data cubes. Inspired by masked image modeling pretraining, we also devise efficient masking strategies for pretraining each of the spectral and spatial transformers. We conduct experiments on six publicly available datasets for HSI classification task and demonstrate that our model achieves state-of-the-art performance in all the datasets. The code for our model will be made available at https://github.com/csiro-robotics/FactoFormer.
@@
Hyperspectral image, transformer network, factorized, self-supervised pretraining.
@@ I Introduction
However, previous methods  [13, 14] mainly applied pure Vision Transformer (ViT) architecture for spectral encoding, using a simple yet effective spectral group tokenization strategy, but they have not fully exploited the spatial and spectral interactions present in hyperspectral image cubes. Furthermore, the limitation of annotated data is a significant challenge when training ViTs, as they rely heavily on large-scale training data. However, most of the available hyperspectral datasets contain limited labeled proportions, such as the Indian Pines, University of Pavia, and Houston 2013 datasets, which possess only 48.75%, 17.43%, and 2.26% of annotated data, respectively. Recently, Masked Image Modeling (MIM) [16, 17] has been introduced as a simple yet powerful self-supervised pretraining framework for Vision Transformers (ViT) without the use of labels. However, the MIM pretraining strategies often have been limited to a single modality (RGB).
@@
The proposed factorized architecture not only increases the efficiency and scalability of modeling spectral and spatial interactions but also enables factorized self-supervised pretraining. To mitigate the lack of large labeled data, we design novel masking strategies to effectively pretrain spectral and spatial transformers individually, leveraging a vast amount of unlabeled hyperspectral data. A random proportion of the input tokens are masked, and only the visible tokens are passed through spectral and spatial transformers, respectively. The latent space representations from each transformer, combined with the masked tokens, are then used to reconstruct the original information for the masked regions.
@@
  • We propose novel self-supervised pretraining strategies for the factorized spatial and spectral transformers, which allows us to better utilize the large proportion of unlabeled data in the public datasets.
@@ II-C Transformers for Hyperspectral Images
Transformer-based networks have revolutionized the entire natural language processing domain with a significant performance boost. Inspired by this, vision transformer (ViT) [27] was introduced later, where images are treated as a sequence of patches. This model achieved state-of-the-art performance for image classification with large-scale pretraining. Transformers have proven to be the ideal architecture for capturing long range dependencies through its multi-head self attention mechanism which processes input as a sequence of tokens. Considering the large number of spectral bands in HSI, which has high correlation and granular information, extracting these global dependencies are crucial for learning better features. Therefore, utilizing transformers and adapting it for learning spatial-spectral features in HSI would lead to better performance. However, there are only few works exploring the transformer architectures for HSI and its full potential in the domain is yet to be explored. Sun et al. [15] proposed a hybrid network that utilizes both CNNs and transformers. Initially, 3D and 2D CNNs are used to extract features, and then the extracted features are tokenized by Gaussian weighted tokenizer and sent to the transformer encoder. In this method, the main feature extraction is based on CNNs, where a single-layer transformer is utilized at the last layer to learn relationships between high-level semantic features. Mei et al. [28] proposed a group-aware hierarchical transformer, which has three stages where each stage is composed of grouped pixel embedding and transformer layers. Here, the proposed grouped pixel embedding comprises convolutions to learn local features, while the transformer is utilized to learn global dependencies. These hybrid approaches perform the initial feature extraction using convolutions where the transformers are used in order to find the global dependency of extracted local features. They did not explore how transformers can be directly used with HSI to learn spectral-spatial feature interactions without convolutions. SpectralFormer [13] utilized the ViT architecture and proposed a pure transformer-based network for HSI with an additional skip connection mechanism to fuse information between transformer layers. Here, the input is patched along the spectra by grouping neighboring bands. Even though the model performed well, the input samples are patched only along the spectral dimension, where the spatial information in HSIs is not fully exploited. Also, the model is trained in a fully supervised manner using limited training samples, which degrades the performance of transformers due to the lack of annotated data. Recently, MAEST [14] utilizes a similar transformer architecture as SpectralFormer[13] and added a self-supervised pretraining framework for only spectral dimension, which is briefly discussed in the next section. Peng et al. [29] proposed a spatial-spectral transformer architecture with cross-attention for HSI. Here, the input is tokenized in spatial and spectral dimensions and fused with a spatial-spectral cross-attention layer in the transformer block. In this method, principal component analysis is used initially to reduce the dimensionality; hence the model is not exposed to full spectral ranges which may cause the loss of fine-grained information. Also, the approach is fully supervised and trained with limited labeled data and the additional convolution layers used prior to tokenization adds more complexity to the model.
@@ II-D Self-Supervised Learning
In this section we have explored and analyzed related works that closely align with our proposed method. We propose FactoFormer: a factorized transformer with self-supervised pretraining to learn rich spectral-spatial feature representations from hyperspectral data. The spectral-spatial factorization is proposed in [29] as well, but the approach is fully supervised. Also, the input is pre-processed with PCA in order to reduce the spectral dimension, potentially leading to a loss of fine-grained information. The input is further processed with additional convolutions prior to sending it to the transformer. In contrast, FactoFormer adopts a different approach. We tokenize the raw input without any preprocessing and directly feed it to the transformers. This approach enables the model to learn more fine-grained spatial-spectral correlations present in the hyperspectral data, making it well-suited for the task at hand.In terms of self-supervised pretraining both MAEST [14] and proposed FactoFormer utilized the unlabelled data to pretrain the network. In MAEST [14], the input is patched only along the spectra by grouping neighbouring bands, whereas in FactoFormer we propose a factorized pretraining approach which enables learning rich spectral and spatial features separately during pretraining and jointly fine-tuning to learn the higher-order spectral-spatial features. Also compared to MAEST [14] our proposed method is computationally more efficient due to the simple linear layer decoder used to reconstruct only masked tokens whereas in MAEST [14] a heavy transformer-based decoder is used to reconstruct the full input.
@@ IV Methodology
In this section, we first give an overview of our proposed method FactoFormer. Next, we discuss the proposed novel factorized transformer architecture. Finally, we explain the self-supervised pretraining strategy used in our method. The overall architecture of the model is shown in Figure 1.
@@ IV-A FactoFormer
Our proposed FactoFormer learns spectral and spatial representations from hyperspectral data using factorized transformers. We process the entire input cube using both spatial and spectral transformers in order to capture the intrinsic spatial-spectral correlations. But the way we tokenize and compute self-attention to learn salient spatial and spectral information discriminates against one another. Further, with the fusion mechanism employed at the end, we capture more fine-grained higher-order spatial-spectral correlations. The factorized architecture enables factorized pretraining, where both transformers can be treated separately and pretrained using specific self-supervised pre-text tasks.
@@
Figure 2: An overview of the proposed pretraining network for spectral transformer in FactoFormer.
IV-B Self-Supervised Pretraining in FactoFormer
Inspired by Masked Image Modeling (MIM), we design a new pretraining strategy for both spectral and spatial transformers. We utilized the large proportion of unlabeled data from the same dataset for self-supervised pretraining. As these unlabeled samples belong to the same image, environmental conditions, instrumental noise, and other affecting factors remain constant during pretraining. They are helpful for learning better feature representations. Both transformers are pretrained separately with spatially consistent and spectrally consistent masking strategies to enhance their underlying feature representations.
Pretraining of Spectral Transformer: We apply patching to the input sample in the spectral dimension, followed by linear embedding and positional encoding as described in Eq. 6. Subsequently, we randomly mask tokens from the input sample, passing only the visible tokens through the spectral transformer. The latent space representations from the spectral transformer, combined with the masked tokens, are then used to reconstruct the original information for the masked regions. For this purpose, we utilize a simple linear layer as the decoder, in contrast to the heavy transformer-based decoder used in MAEST [14]. Finally, we use Mean Squared Error (MSE) as the loss function. As shown in Figure 2, masks are spatially consistent. Each token corresponds to the complete spatial information of a particular band in the input sample, allowing the model to learn better representations along the spectral dimension from spectrally patched tokens.
Pretraining of Spatial Transformer: We apply patching to the input in the spatial dimension and pass them through the spatial transformer along with positional embeddings as outlined in Eq. 8. A random proportion of the input tokens are masked, and only the visible tokens are passed through the spatial transformer. Similar to the spectral transformer, we utilize a simple linear layer as the decoder to reconstruct the masked regions and use MSE as the loss function. Notably, the spectrally consistent masks enable the transformer to learn meaningful representations from spatially patched tokens, as depicted in Figure 3.
Figure 3: An overview of the proposed pretraining network for spatial transformer in FactoFormer.
End-to-End Fine-tuning: We use the pretrained weights of the spectral and spatial transformers to initialize FactoFormer. Then we fine-tune the network end-to-end in order to perform classification. Our proposed factorization technique enables FactoFormer to exploit the pretrained weights and learn rich spectral and spatial feature representations jointly without imposing significant complexity to the network.
@@ V-A Datasets
We use six popular hyperspectral datasets to evaluate the performance of the proposed model and to compare it with other state-of-the-art methods. These datasets include Indian Pines [34], University of Pavia [35], and Houston 2013 [36], as well as the Wuhan UAV-borne hyperspectral image (WHU-Hi) dataset [37], which consists of three individual datasets: WHU-Hi-LongKou, WHU-Hi-HanChuan, and WHU-Hi-HongHu. The details on number of samples used for pretraining, fine-tuning and testing are shown in Table I.
@@
TABLE I: Details of Data used for pretraining, Fine-tuning and Testing
@@ V-B Implementation Details
Network Architectures: We use a transformer architecture with five layers and four attention heads for both spectral and spatial transformers.We set the embedding size for spectral and spatial transformers as 32 and 64, respectively. Table II outlines the details of the network architectures. The same architectures were preserved during pretraining as well.
@@
Pretraining: We employ the Adam optimizer with a linear learning rate scheduler and train for 200 epochs. The batch size is set to 32 with an initial learning rate of $5e^{-4}$ and decayed by a factor of 0.9 for every 20 epochs. The pretraining setting is detailed in Table IV.
Fine-tuning: For fine-tuning, the pretrained weights are used to initialize both spectral and spatial transformers and fine-tuned end-to-end. A learning rate search is performed, and the best for each dataset is reported. We use the Adam optimizer with a linear learning rate scheduler and set the batch size to 32. For both the Indian Pines and University of Pavia datasets, the fine-tuning model converges after 80 epochs, while for Houston and WHU-Hi datasets, it converges in just 40 epochs. The Finetuning setup is detailed in Table IV.
TABLE III: Pretraining Setting
@@ V-C State-of-the-art Methods
Transformer Based Networks: For transformer-based networks, we used the ViT architecture [27], SpectralFormer [13], and MAEST [14]. The ViT uses five layers, four attention heads, an embedding size of 64, and an MLP with a hidden dimension of 8. SpectralFormer[13] uses the same settings as the ViT with cross-layer adaptive fusion (CAF). CAF employs a skip connection mechanism to integrate the features of skipping encoder blocks. MAEST[14] is a self-supervised approach that utilizes Masked Autoencoders (MAE) to train an encoder-decoder network that reconstructs masked patches. Then a classification encoder with similar network architecture as SpectralFormer[13] is initialized with the pretrained encoder weights and fine-tuned to perform classification.
@@ VI-A Comparison with the State-of-the-Art
When comparing FactoFormer with MAEST[14], which is the current state-of-the-art method using a transformer with a self-supervised pretraining approach, our method outperforms it with 7.15%, 4.13%, and 0.58% higher OA on Indian Pines, University of Pavia and Houston 2013 datasets respectively. It shows that our novel factorized approach with self-supervised pretraining improves performance significantly. Also, interestingly our models converge with less than 100 epochs during fine-tuning on all three datasets, whereas, in MAEST[14], the models are fine-tuned for more than 200 epochs. It again highlights that the factorized pretraining in FactoFormer learns rich feature representations, which enables fast convergence during fine-tuning. Also, FactoFormer is computationally more efficient in terms of the number of model parameters and run-time, which is briefly discussed in Section VI-C under computational efficiency.
Compared to SpectralFormer [13], which is a fully supervised transformer-based approach, FactoFormer outperforms it by 9.54% on Indian Pines, 4.12% on University of Pavia and 1.12% on Houston 2013 datasets highlighting the importance of factorization and pretraining. ViT[27] shows the least performance of all the transformer methods, as it is applied directly as in images for spectral tokens in HSI without adapting it to suit the structure of hyperspectral data. In general, SpectralFormer[13], MAEST[14], and FactoFormer perform better than all other classic backbone networks and conventional classifiers in all three datasets, demonstrating the effectiveness of transformer-based approaches for HSI. 2-D CNN[9] performs better compared to 1-D CNN[8], RNN[19] and miniGCN[41] on all three datasets. When comparing 2-D CNN [9], FactoFormer achieves 15.41%, 9.14%, and 5.41% higher accuracy on Indian Pines, University of Pavia and Houston 2013 datasets, respectively. This significant performance gap exhibits the effectiveness of FactoFormer over CNN-based approaches.